Coolbrook Company has the following information available for the past year:   

River Division Stream Division
Sales revenue $ 1,218,000 $ 1,803,000
Cost of goods sold and operating expenses 887,000 1,297,000
Net operating income $ 331,000 $ 506,000
Average invested assets $ 1,030,000 $ 1,510,000

   

The company’s hurdle rate is 7.76 percent.

Required:

1. Calculate return on investment (ROI) and residual income for each division for last year.

River Division Stream Division
ROI % %
Residual Income (Loss)

2. Recalculate ROI and residual income for the division for each independent situation that follows:

a. Operating income increases by 12 percent.

River Division Stream Division
ROI % %
Residual Income (Loss)

b. Operating income decreases by 10 percent.

River Division Stream Division
ROI % %
Residual Income (Loss)

c. The company invests $250,000 in each division, an amount that generates $109,000 additional income per division.

River Division Stream Division
ROI % %
Residual Income (Loss)

d. Coolbrook changes its hurdle rate to 5.76 percent.

River Division Stream Division
ROI % %
Residual Income (Loss)
Answer #1

(1)

River Division Stream Division
ROI 32.14% 33.51%
Residual Income $251072 $388824

Explanations :-

ROI = Net operating Income/Average Invested assets

River Division = $331000/$1030000 = 32.14%

Stream Division = $506000/$1510000 = 33.51%

Residual Income = Net operating income - (Average Invested Assets * Hurdle Rate)

River Division = $331000 - ($1030000 * 7.76%) = $251072

Stream Division = $506000 - ($1510000 * 7.76%) = $388824

(2a)

River Division Stream Division
ROI 35.99% 37.53%
Residual Income $290792 $449544

Explanations :-

ROI = Net operating Income/Average Invested assets

River Division = ($331000 * 112%) / $1030000 = 35.99%

Stream Division = ($506000 * 112%) / $1510000 = 37.53%

Residual Income = Net operating income - (Average Invested Assets * Hurdle Rate)

River Division = ($331000 * 112%) - ($1030000 * 7.76%) = $290792

Stream Division = ($506000 * 112%) - ($1510000 * 7.76%) = $449544

(2b)

River Division Stream Division
ROI 35.99% 37.53%
Residual Income $217972 $338224

Explanations :-

ROI = Net operating Income/Average Invested assets

River Division = ($331000 * 90%) / $1030000 = 28.92%

Stream Division = ($506000 * 90%) / $1510000 = 30.16%

Residual Income = Net operating income - (Average Invested Assets * Hurdle Rate)

River Division = ($331000 * 90%) - ($1030000 * 7.76%) = $217972

Stream Division = ($506000 * 90%) - ($1510000 * 7.76%) = $338224

(2c)

River Division Stream Division
ROI 34.38% 34.94%
Residual Income $340672 $478424

Explanations :-

ROI = Net operating Income/Average Invested assets

River Division = ($331000 + $109000) / ($1030000 + $250000) = 34.38%

Stream Division = ($506000 + $109000) / ($1510000 + $250000)= 34.94%

Residual Income = Net operating income - (Average Invested Assets * Hurdle Rate)

River Division = ($331000 + $109000) - ($1280000 * 7.76%) = $340672

Stream Division = ($506000 + $109000) - ($1760000 * 7.76%) = $478424

(2d)

River Division Stream Division
ROI 32.14% 33.51%
Residual Income $271672 $419024

Explanations :-

ROI = Net operating Income/Average Invested assets

River Division = $331000/$1030000 = 32.14%

Stream Division = $506000/$1510000 = 33.51%

Residual Income = Net operating income - (Average Invested Assets * Hurdle Rate)

River Division = $331000 - ($1030000 * 5.76%) = $271672

Stream Division = $506000 - ($1510000 * 5.76%) = $419024